Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: Recovery from power outage, spate of new software updates. Resumed science observing (with limitations) Saturday afternoon. Issues with filter gain settings caused issues with narrow band observations through Monday morning. Scheduling done via OST exclusively. Double maintenance day this week (Tues-Wed). Thursday morning discussion for spectral line group is postponed due to the HR benefits/open enrollment meeting at the same time (Momjian on Zeeman testing results - next week hopefully).
  • Key dates:
  • Staff Travel:
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD

Hardware status and updates (Langley, Perley, Rupen)

Software status and updates (Butler)

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

Double maintenance day week:
  • Resolution of Brent's visit. Done.
  • Resolution of narrow band observing restrictions. Done.
  • Science (planned Thursday-Friday day time observing sessions)
  • Other testing/issue resolutions.
  • Interrupt status for additional 3-bit testing

Commissioning Reports

Solar: Perley: Results of solar mode observing discussions with Tim

  • Also CCB; L band reverse coupler scheme (supporting full range of solar phenomena is moving ahead); Tim+community have also asked for support to upgrade the SCXKa with 20 dB coupler and solar cals (rather than just support the Ka band observing); this supports the active sun up to solar flares. Funding will need to wait until resolution of the 3-bit planning is completed. Plan for solar tests this week (initial completed). Note: Rick will also discuss travels at Science Tea today.

Calibration: Greisen/Dhawan: Switched power testing (follow-up)

  • AIPS is working with the switched power data; there are a suite of utilities:
    • SNPLT: plots SY table
    • SNEDT: edits uv-data, SY table and allows applying uv flags (for off source data).
    • EDITA: edit uv-data based on SY table
    • TYSMO: smooth/clip TY, SY tables
    • TYAPL: undo/do application of a particular SY table
    • Application of these leads to a variation of the average amplitude of .98-.99 (versus 5-10 normally). New information on this is available in the Appendix E.
  • 1dB on/off tests
    • Noted that the first 1-2 records are not correct (jumpy values); need to be flagged. Otherwise this test performed as expected.
    • Note the identical values for different subbands issue (reported last week) has now been resolved.

Calibration: Momjian: Receiver SEFD monitoring project commencement.

  • Started: seeing some bad values on particular receivers (under investigation). Overall initial baseline for data is set though the methodology will evolve.

Array: Dhawan/Rupen: RF switch tests with rapid monitoring

  • Data taken but not yet analyzed.

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Correlator: Krauss/Hunter: 7 vs 8 node observing (follow-up; pending RSRO results)

Commissioning Test Planning

  • OPT: webtest of 1.06 capabilities/bug fixes (*); Deployed today.
  • Correlator: CBE averaging with restricted nodes
  • Calibration: Switched power tracking over several airmasses.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Scheduled for double maintenance day this week. Postponed until wideband X available on EA24
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Spectral Line: Momjian: Zeeman test follow-upHigh frequency/low frequency results.


  • Sent a clarification on key information to include; comments?

10B-154: Owen (first results)

  • Test data taken. Noted difficulties given the delay in scheduling which have shifted the project into the day time and required tracking ionosphere issues. Exploring polarization calibration - polarization will be meeting this week but key is understanding which calibrators (e.g., not 3C48) can be useful at which lower frequencies. More soon.

Other updates:

  • Adam - posted updates to (things are looking good!).
  • Greg - ran a test block with 20 min scans; was able to fill with OBIT so the data were salvaged and usable; still pending the green light to go ahead with program.

  • Notes:
    • Current Residents
      • Satoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • Hauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • Greg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
      • Adam Leroy (POC: TBD); area(s):; 10A-196 (AL746): Resolving the Starbursts in Nearby LIRGs and ULIRGs
      • Jill Knapp (POC: Rupen/Claussen); area(s): recirculation; IRC+10216 demo science; 10B-200;10B-186. Welcome again.

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Commissioning Group updates


  • Frazer noted that the OPT/OST does not readily allow restrictions based on proximity to the sun. In addition, the current OST weather trending (phase for 10 minutes, wind for 2 hours) used for determining viability needs work. This will be kicked to the scheduling group for discussion/resolution. Some could be requested for development of 1.07.

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.

-- JosephMcMullin - 2010-10-25
Topic revision: r3 - 2010-10-27, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback